Use the calculator provided to solve the following problems.
Consider a t distribution with 8 degrees of freedom. Compute P(-1.92 < t < 1.92). Round your answer to at least three decimal places.
Consider a t distribution with 30 degrees of freedom. Find the value of c such that P(t >_ c) = 0.01. Round your answer to at least three decimal places.

P(-1.92 < t < 1.92) = ____

c = ____
Answer #1

Given degrees of freedom is 8 and P(-1.92 < t < 1.92) = 0.9089

since P(-1.92 < t < 1.92) = P(t<1.92) - P(t < -192) = 0.9544 - 0.04556 =0.9089

also

and

Given P(t > = c) = 0.01 the value of C is 2.457

since